# Residue-Enhanced Pion–Rho Mixing as the Origin of Nonmonotonic Charged Pion Mass in Magnetic Fields

### v1

@@ Abstract
Effective near-pole description. To treat $\pi^{+}$ and $\rho^{+}$ as quark–antiquark bound states on equal footing and to describe their magnetic-field–induced mixing consistently, we derive the quadratic bosonic action from a microscopic quark theory in a uniform external magnetic field. Projecting the resulting inverse propagator onto mesonic Landau eigenstates yields a Landau-level–resolved kernel that contains the full energy dependence inherited from the quark loop.
@@
where $E_{\pi}(B)$ and $E_{\rho}(B)$ are are the unmixed lowest-Landau-level energies of the pion and longitudinal rho meson, respectively. The physical eigenmodes within this near-pole effective theory are obtained from the pole condition
@@
This procedure assumes that the relevant collective mode remains adiabatically connected to the unmixed lowest-Landau-level pole, an assumption whose validity can be tested by comparison with the full Landau-projected kernel introduced below.
@@
The structure of the kernel makes the physical mechanism transparent: as the magnetic field increases, the mixing induces level repulsion between a light pion mode $\pi^{+}$ and a heavier longitudinal rho mode $\rho^{+}_{s_{z}=0}$. When the residue $Z(B)$ decreases rapidly, the effective mixing strength $h(B)$ is strongly amplified near the pole, leading to a downward bending of the lower eigenmode. This residue-enhanced level repulsion provides a natural route to non-monotonic magnetic-field dependence of the charged pion energy.
Microscopic origin We now derive the ingredients entering the near-pole effective kernel directly from the underlying quark dynamics in a magnetic field. In particular, the unmixed energies $E_{\pi}(B)$, $E_{\rho}(B)$, the wave-function renormalizations $Z_{\pi}(B)$, $Z_{\rho}(B)$, and the loop-induced mixing $g_{\rho\pi}^{\rm loop}(B)$ are obtained from Landau-level–projected quark polarization functions. We evaluate these quantities within the SU(2) Nambu–Jona-Lasinio model with quarks minimally coupled to the background field $\mathcal{L}=\bar{\psi}(iD\!\!\!\!/-m_{0})\psi+G_{S}[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}\psi)^{2}]-G_{V}[(\bar{\psi}\gamma^{\mu}\tau^{a}\psi)^{2}]$, with $\tau^{a}$ representing the isospin Pauli matrices, and $G_{S}$ and $G_{V}$ the coupling constants with respect to the scalar (pseudoscalar) and the vector channels Klevansky (1992); Hatsuda and Kunihiro (1994); Vogl and Weise (1991). The covariant derivative couples quarks to an external magnetic field $\vec{B}=B\hat{z}$ along the positive $z$ direction. For the quark propagator we use Landau-level representations Schwinger (1951); Miransky and Shovkovy (2015).
@@
where we restrict to the reduced subspace of $\pi^{+}$ and $\rho^{+}_{s_{z}=0}$ denoted by $\Phi=(\pi,\rho_{3})^{T}$, $\mathcal{G}^{-1}=\text{diag}(1/2G_{S},1/2G_{V})$ is the contact inverse propagator, and $\Pi$ is the polarization matrix from the quark loop. $S^{(2)}_{\rm tree}$ comes from the gauge-invariant tree-level counterterm
@@
with $\phi$ stands for $\pi^{+}$ or the longitudinal $\rho^{+}$, and $\varphi_{n}$ is the standard gauge dependent Landau wavefunctions. Because u and d quarks carry different electric charges, the polarization functions acquire nontrivial Schwinger phases. Upon projection onto charged-meson Landau eigenstates, these phases combine with the external wave functions to produce gauge-invariant Landau-projected kernels. Landau quantization therefore enters both through internal quark propagators and external meson legs, and their combined structure determines the magnetic-field dependence of the meson spectrum. Then the boson quadratic action (7) becomes a sum over Landau levels:
@@
with $\Phi_{n}=(\pi_{n},\rho_{n})^{T}$ is the subspace of $\pi^{+}$ and $\rho^{+}_{s_{z}=0}$ at given Landau level $n$. Projecting the quadratic bosonic action onto mesonic Landau eigenstates yields the microscopic Landau-level–projected inverse propagator,
@@
with $\mathcal{K}_{\pi\pi}=\frac{1}{2G_{S}}-\Pi_{\pi\pi,n}(q_{0},B)$, $\mathcal{K}_{\rho\rho}=\frac{1}{2G_{V}}-\Pi_{\rho\rho,n}(q_{0},B)$, $\mathcal{K}_{\rho\pi}=-{\Pi}^{3}_{\rho\pi,n}(q_{0},B)+\mathcal{K}_{\rho\pi}^{\text{tree}}(q_{0},B)$, $\mathcal{K}_{\pi\rho}=-{\Pi}^{3}_{\pi\rho,n}(q_{0},B)+\mathcal{K}_{\pi\rho}^{\text{tree}}(q_{0},B)$, and the Landau projected polarization function is defined as $\Pi_{\phi_{1}\phi_{2},n}(q_{0},B)=\int d^{4}xd^{4}y\varphi_{n}^{*}(x)\Pi_{\phi_{1}\phi_{2}}(x,y;B)\varphi_{n}(y)$. This Landau-level–projected kernel contains the full energy dependence inherited from the quark loop and does not rely on any near-pole expansion. Solving $\text{det}\mathcal{K}(q_{0};N,n=0)=0$ directly therefore yields the complete determinant spectrum within the NJL model, providing a reference against which the near-pole reduction can be assessed.
The unmixed lowest-Landau-level energies $E_{\pi}(B)$ and $E_{\rho}(B)$ are defined by the pole conditions of the diagonal Landau-projected kernels,
@@
We first determine the unmixed lowest-Landau-level (LLL) energies of the charged pion and longitudinal rho meson from the diagonal Landau-projected kernels. As shown in Fig.1, both $E_{\pi}(B)$ and $E_{\rho}(B)$ increase monotonically with magnetic field and exhibit the expected $\sqrt{eB}$ behavior at large fields, reflecting Landau quantization. When converted to rest masses via $m_{\phi}=\sqrt{E_{\phi}^{2}-eB}$, the charged pion mass remains monotonic. The strict monotonicity of the unmixed LLL energies demonstrates that Landau quantization and magnetic catalysis alone cannot generate a turnover.
@@
Solving the pole condition (4) yields two mixed eigenmodes. The lower eigenmode $E_{-}$ is shown in Fig.4. At weak fields it closely follows the unmixed pion LLL energy. At intermediate fields, increasing mixing induces level repulsion between the light pion mode and the heavier longitudinal rho mode, causing the lower eigenvalue to bend downward. This turnover is a direct consequence of residue-enhanced pion–rho mixing within the same Landau level. For comparison, Fig.4 also shows the result obtained by solving $\text{det}\,\mathcal{K}(q_{0})=0$ directly using the full Landau-level–projected kernel (13). At weak and intermediate fields the two approaches agree, confirming that the near-pole formulation captures the essential determinant physics. At larger $eB$ the increasing deviation reflects sensitivity to higher-order energy dependence and provides a quantitative estimate of systematic uncertainties in the truncated near-pole description.
@@
While the precise location of the maximum depends on quantitative details such as the magnetic-field dependence of $Z_{\rho}(B)$ and the effective pion–rho coupling, the existence of the turnover does not rely on fine tuning. In the near-pole framework, it is controlled by the competition between Landau-level separation and residue-enhanced mixing. Corrections beyond the present treatment—such as nonlocal interactions or continuum effects in the polarization functions—may shift the peak position without altering the underlying mechanism.
@@
Summary and Discussion In this Letter, we have identified a clean microscopic mechanism underlying the non-monotonic magnetic-field dependence of the charged pion mass observed in lattice QCD. Starting from a quark-level description and formulating the problem in terms of Landau-projected near-pole dynamics, we showed that the lowest-Landau-level charged pion mixes with the longitudinally polarized charged rho meson in a magnetic background. This symmetry-allowed mixing induces level repulsion, which is dynamically amplified by a rapid suppression of the rho-meson wave-function renormalization near the pole. The resulting residue-enhanced mixing drives the downward bending of the lower eigenmode at strong magnetic fields.
A central implication of this analysis is that, in the presence of strong mixing, lattice observables are governed by the lowest near-pole eigenmode rather than by unmixed Landau levels or rest-mass shifts alone. Residue effects—typically subleading in vacuum spectroscopy—become decisive in magnetic backgrounds.

### v2

@@ Abstract
Effective near-pole description. To treat $\pi^{+}$ and $\rho^{+}$ as quark-antiquark bound states on equal footing and to describe their magnetic field induced mixing consistently, we derive the quadratic bosonic action from a microscopic quark theory in a uniform external magnetic field. Projecting the resulting inverse propagator onto mesonic Landau eigenstates yields a Landau-level-resolved kernel that contains the full energy dependence inherited from the quark loop.
@@
where $E_{\pi}(B)$ and $E_{\rho}(B)$ are are the unmixed lowest Landau level energies of the pion and longitudinal rho meson, respectively. The physical eigenmodes within this near-pole effective theory are obtained from the pole condition
@@
This procedure assumes that the relevant collective mode remains adiabatically connected to the unmixed lowest Landau level pole, an assumption whose validity can be tested by comparison with the full Landau-projected kernel introduced below.
@@
The structure of the kernel makes the physical mechanism transparent: as the magnetic field increases, the mixing induces level repulsion between a light pion mode $\pi^{+}$ and a heavier longitudinal rho mode $\rho^{+}_{s_{z}=0}$. When the residue $Z(B)$ decreases rapidly, the effective mixing strength $h(B)$ is strongly amplified near the pole, leading to a downward bending of the lower eigenmode. This residue enhanced level repulsion provides a natural route to non-monotonic magnetic-field dependence of the charged pion energy.
Microscopic origin We now derive the ingredients entering the near-pole effective kernel directly from the underlying quark dynamics in a magnetic field. In particular, the unmixed energies $E_{\pi}(B)$, $E_{\rho}(B)$, the wave-function renormalizations $Z_{\pi}(B)$, $Z_{\rho}(B)$, and the loop induced mixing $g_{\rho\pi}^{\rm loop}(B)$ are obtained from Landau level projected quark polarization functions. We evaluate these quantities within the SU(2) Nambu–Jona-Lasinio model with quarks minimally coupled to the background field $\mathcal{L}=\bar{\psi}(iD\!\!\!\!/-m_{0})\psi+G_{S}[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}\psi)^{2}]-G_{V}[(\bar{\psi}\gamma^{\mu}\tau^{a}\psi)^{2}]$, with $\tau^{a}$ representing the isospin Pauli matrices, and $G_{S}$ and $G_{V}$ the coupling constants with respect to the scalar (pseudoscalar) and the vector channels Klevansky (1992); Hatsuda and Kunihiro (1994); Vogl and Weise (1991). The covariant derivative couples quarks to an external magnetic field $\vec{B}=B\hat{z}$ along the positive $z$ direction. For the quark propagator we use Landau-level representations Schwinger (1951); Miransky and Shovkovy (2015).
@@
where we restrict to the reduced subspace of $\pi^{+}$ and $\rho^{+}_{s_{z}=0}$ denoted by $\Phi=(\pi,\rho_{3})^{T}$, $\mathcal{G}^{-1}=\text{diag}(1/2G_{S},1/2G_{V})$ is the contact inverse propagator, and $\Pi$ is the polarization matrix from the quark loop. $S^{(2)}_{\rm tree}$ comes from the gauge invariant tree-level counterterm
@@
with $\phi$ stands for $\pi^{+}$ or the longitudinal $\rho^{+}$, and $\varphi_{n}$ is the standard gauge dependent Landau wavefunctions. Because u and d quarks carry different electric charges, the polarization functions acquire nontrivial Schwinger phases. Upon projection onto charged meson Landau eigenstates, these phases combine with the external wave functions to produce gauge invariant Landau-projected kernels. Landau quantization therefore enters both through internal quark propagators and external meson legs, and their combined structure determines the magnetic field dependence of the meson spectrum. Then the boson quadratic action (7) becomes a sum over Landau levels:
@@
with $\Phi_{n}=(\pi_{n},\rho_{n})^{T}$ is the subspace of $\pi^{+}$ and $\rho^{+}_{s_{z}=0}$ at given Landau level $n$. Projecting the quadratic bosonic action onto mesonic Landau eigenstates yields the microscopic Landau level projected inverse propagator,
@@
with $\mathcal{K}_{\pi\pi}=\frac{1}{2G_{S}}-\Pi_{\pi\pi,n}(q_{0},B)$, $\mathcal{K}_{\rho\rho}=\frac{1}{2G_{V}}-\Pi_{\rho\rho,n}(q_{0},B)$, $\mathcal{K}_{\rho\pi}=-{\Pi}^{3}_{\rho\pi,n}(q_{0},B)+\mathcal{K}_{\rho\pi}^{\text{tree}}(q_{0},B)$, $\mathcal{K}_{\pi\rho}=-{\Pi}^{3}_{\pi\rho,n}(q_{0},B)+\mathcal{K}_{\pi\rho}^{\text{tree}}(q_{0},B)$, and the Landau projected polarization function is defined as $\Pi_{\phi_{1}\phi_{2},n}(q_{0},B)=\int d^{4}xd^{4}y\varphi_{n}^{*}(x)\Pi_{\phi_{1}\phi_{2}}(x,y;B)\varphi_{n}(y)$. This Landau level projected kernel contains the full energy dependence inherited from the quark loop and does not rely on any near-pole expansion. Solving $\text{det}\,\mathcal{K}(q_{0};N,n=0)=0$ directly therefore yields the complete determinant spectrum within the NJL model, providing a reference against which the near-pole reduction can be assessed.
The unmixed LLL energies $E_{\pi}(B)$ and $E_{\rho}(B)$ are defined by the pole conditions of the diagonal Landau-projected kernels,
@@
We first determine the unmixed LLL energies of the charged pion and longitudinal rho meson from the diagonal Landau-projected kernels. As shown in Fig.1, both $E_{\pi}(B)$ and $E_{\rho}(B)$ increase monotonically with magnetic field and exhibit the expected $\sqrt{eB}$ behavior at large fields, reflecting Landau quantization. When converted to rest masses via $m_{\phi}=(E_{\phi}^{2}-eB)^{1/2}$, the charged pion mass remains monotonic. The strict monotonicity of the unmixed LLL energies demonstrates that Landau quantization and magnetic catalysis alone cannot generate a turnover.
@@
Solving the pole condition (4) yields two mixed eigenmodes. The lower eigenmode $E_{-}$ is shown by the red line in Fig.4. At weak fields it closely follows the unmixed pion LLL energy. At intermediate fields, increasing mixing induces level repulsion between the light pion mode and the heavier longitudinal rho mode, causing the lower eigenvalue to bend downward. This turnover is a direct consequence of residue-enhanced pion-rho mixing within the same Landau level. For comparison, in Fig.4 the black line shows the result obtained by solving $\text{det}\,\mathcal{K}(q_{0})=0$ directly using the full Landau level projected kernel (13). At weak and intermediate fields the two approaches agree, confirming that the near-pole formulation captures the essential determinant physics. At larger $eB$ the increasing deviation reflects sensitivity to higher-order energy dependence and provides a quantitative estimate of systematic uncertainties in the truncated near-pole description.
@@
While the precise location of the maximum depends on quantitative details such as the magnetic field dependence of $Z_{\rho}(B)$ and the effective pion-rho coupling, the existence of the turnover does not rely on fine tuning. In the near-pole framework, it is controlled by the competition between energy separation and residue-enhanced mixing. Corrections beyond the present treatment—such as nonlocal interactions or continuum effects in the polarization functions—may shift the peak position without altering the underlying mechanism.
@@
Summary and Discussion In this Letter, we have identified a clean microscopic mechanism underlying the non-monotonic magnetic field dependence of the charged pion mass observed in lattice QCD. Starting from a quark-level description and formulating the problem in terms of Landau-projected near-pole dynamics, we showed that the lowest Landau level charged pion mixes with the longitudinally polarized charged rho meson in a magnetic background. This symmetry allowed mixing induces level repulsion, which is dynamically amplified by a rapid suppression of the rho meson wave-function renormalization near the pole. The resulting residue-enhanced mixing drives the downward bending of the lower eigenmode at strong magnetic fields.
A central implication of this analysis is that, in the presence of strong mixing, lattice observables are governed by the lowest near-pole eigenmode rather than by unmixed Landau levels or rest-mass alone. Residue effects—typically subleading in vacuum spectroscopy—become decisive in magnetic backgrounds.